# Transformers as Unsupervised Learning Algorithms: A study on Gaussian Mixtures

### v1

@@ C.1 Formal statement of Theorem 1
where $\overline{X}_{i}=[{X}_{i}^{\top},{\mathbf{0}}_{d_{0}-d}^{\top}]^{\top}$, $\overline{{\boldsymbol{\pi}}}_{\log}=[{\boldsymbol{\pi}}_{\log}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{\mu}_{i\%K}=[\mu_{i\%K}^{\top}$, ${\mathbf{0}}_{d_{0}-d}^{\top}]^{\top}$, $\overline{{\mathbf{w}}}_{i}=[{\mathbf{w}}_{i}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{{\mathbf{w}}}_{i\log}=[{\mathbf{w}}_{i\log}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{{\boldsymbol{\pi}}}=[{\boldsymbol{\pi}}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $c_{i\%K}\in\mathbb{R}$ and ${\mathbf{e}}_{i\%K}\in\mathbb{R}^{K_{0}}$ denotes the ${i\%K}$-th standard unit vector. For the initialization, we choose ${\boldsymbol{\pi}}_{\log}=\log{\boldsymbol{\pi}}^{(0)}$, $\mu_{i}=\mu_{i}^{(0)}$, $c_{i}=\|\mu_{i}^{(0)}\|_{2}^{2}$. and ${{\boldsymbol{\pi}}}={\mathbf{w}}_{i}={\mathbf{w}}_{i\log}={\mathbf{0}}_{K}$, $i\in[K]$. Finally, take ${\mathbf{H}}^{(0)}={\mathbf{H}}$ which is defined in Eq. 7.
@@
$\widetilde{{\mathbf{h}}}_{i}^{(1)}={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}({\mathbf{H}}^{(0)})\right]}_{:,i}$ $={\mathbf{h}}_{i}^{(0)}+\sum_{j=1}^{N}{\left[\text{SoftMax}{\left({\left({\left({\mathbf{Q}}^{(1)}{\mathbf{h}}_{i}^{(0)}\right)}^{\top}{\left({\mathbf{K}}^{(1)}{\mathbf{h}}_{j}^{(0)}\right)}\right)}_{j=1}^{N}\right)}\right]}_{j}\cdot{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{(0)}$
@@
where ${\overline{\hat{{\mathbf{w}}}}}_{i\log}^{(1)}=\widehat{\log}{\overline{\hat{{\mathbf{w}}}}}_{i}^{(1)}$. Notice that although $\log x$ is not defined at 00, the MLP approximation is well defined with some value which we do not care because we will not use it in the M-step. Similarly, for any %$\ell\%2=1$, $\ell\in\mathbb{N}_{+}$ we have
@@
$\widetilde{{\mathbf{h}}}_{j}^{(2)}={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(2)}}({\mathbf{H}}^{(1)})\right]}_{:,j}$ $={\mathbf{h}}_{j}^{(1)}+\sum_{m=1}^{2}\sum_{i=1}^{N}{\left[\text{SoftMax}{\left({\left({\left({\mathbf{Q}}_{m}^{(2)}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}_{m}^{(2)}{\mathbf{h}}_{j}^{(1)}\right)}\right)}_{j=1}^{N}\right)}\right]}_{j}\cdot{\mathbf{V}}_{m}^{(2)}{\mathbf{h}}_{i}^{(1)}$
@@
Similarly, for any %$\ell\%2=0$, $\ell\in\mathbb{N}_{+}$ we have
@@
Thus, we can get $\hat{{\boldsymbol{\pi}}}^{(\ell)}$ and $\hat{\mu}_{j}^{(\ell)}$, $j\in[K]$ after $2\ell$ layers of transformer constructed above. (The last-layer MLP block retains ${\boldsymbol{\pi}}$ as an output parameter without cleaning it.) Our transformer construction is summarized in Fig. 9, which is the formal version of Fig. 7 in Section 4.2.
Remark C.3.
@@
C.3 Convergence results for EM algorithm
C.3.1 Convergence results for population-EM algorithm
First, we review some notations. Recall that $\pi_{min}=\min_{i}\pi_{i}^{*}$, $\rho_{\pi}=\max_{i}\pi_{i}^{*}/\min_{i}\pi_{i}^{*}$, $R_{ij}=\|\mu_{i}^{*}-\mu_{j}^{*}\|$, $R_{min}=\min_{i\neq j}R_{ij}$ and $R_{max}={\left(\max_{i\neq j}R_{ij}\right)}\vee{\left(\max_{i\in[K]}\left\|{\mu_{i}^{\ast}}\right\|\right)}$. Without the loss of generality, we assume that $R_{max}\geq 1$. For clarity, we restate assumption ((A1)), which is consistent with [24].
@@ C.3.1 Convergence results for population-EM algorithm
The following results gives linear convergenve guarantees of population-EM, which comes from [24].
Theorem C.2 ([24], Theorem 1, part i).
@@
Theorem C.3 ([24], Theorem 1, part ii).
Let $C\geq 128$ be a large enough universal constant. Fix $10<a<$ 1. Suppose the separation condition Eq. 8 holds and suppose the initialization parameter satisfies $D_{m}\leq a$, then $D_{m}^{+}\leq\beta D_{m}$ for some $10<\beta<$ 1.
Remark C.4.
Here the contraction parameter $\beta$ is only dependent with $C$ and $a$. In other words, if we fix $a\in(0,1)$, then for any $\beta\in(0,1)$, there exists a large enough $C$ such that Theorem C.3 holds. For details, see Appendix E in [24].
Combing Theorem C.2 and Theorem C.3, we can get the linear convergence of population-EM algorithm.
C.3.2 Convergence results for empirical-EM algorithm
Now we consider the empirical-EM, i.e., Algorithm A.1. For convenience, the algorithm can be presented as
@@ C.3.2 Convergence results for empirical-EM algorithm
Theorem C.4.
Fix $10<\delta,\beta<$ 1 and $10<a<$ 1. Let $C\geq 128$ be a large enough universal constant. Suppose the separation condition Eq. 8 holds and suppose the initialization parameter satisfies $D_{m}\leq a$. If $n$ is suffcient large such that
@@
where $(i)$ follows from ${|\pi_{i}^{+}-\pi_{i}^{*}|}/{\pi_{i}^{*}}\leq a$. And by Lemma C.3, we have
@@
In conclusion, if we take $\tilde{\beta}=(1-a)/(1+a)\beta$, $\tilde{c}=\tilde{c}_{1}\vee\tilde{c}_{2}$, $C=C(\beta,a)\geq 128$ large enough such that Theorem C.3 holds, and take $\delta_{1}=\delta_{2}=\delta/2$ and use union bound argument, then we get $D_{m}^{+}\leq\beta D_{m}+\varepsilon_{unif}\leq a$.
@@
Lemma C.1 ([40], Lemma B.2.).
Fix $10<\delta<$ 1. Let $B_{1},\ldots,B_{K}\subset\mathbb{R}^{d}$ be Euclidean balls of radii $r_{1},\ldots,r_{K}$. Define $\mathcal{B}=\otimes_{k=1}^{K}B_{k}\subset\mathbb{R}^{Kd}$ and $r=\max_{k\in[K]}r_{k}$. Let $X$ be a random vector in $\mathbb{R}^{d}$ and $W:\mathbb{R}^{d}\times{\cal B}\to\mathbb{R}^{k}$ where $k\leq d$. Assume the following hold:
@@
Remark C.5.
There is one difference between Lemma C.1 and LemmaB.2. in [40]: in Lemma C.1, we use $1+\frac{12nLr}{\delta}$ to replace $\frac{18nLr}{\delta}$, thus we avoid the condition $r_{1},\cdots,r_{K}\geq 1$.
Hence we can get the uniform convergence of $w_{i}(X,\theta)$ and $w_{i}(X,\theta)(X-\mu_{i}^{*})$, $i\in[K]$. Our proof is similar to [40], except that we consider the variation of both $\pi$ and $\mu$. From now on, we denote $\theta_{i}=\{\pi_{i},\mu_{i}\}$, $\theta=\{\theta_{i}\}_{i=1}^{n}$.
Lemma C.2.
[…]
@@
The proof is similar to the proof of Lemma 5.1 in [40]. For simplicity, we omit it. ∎
Lemma C.3.
[…]
@@
The proof is similar to the proof of Lemma 5.4 in [40](Notice that the condition (36) in [40] is trivial in our case). For simplicity, we omit it. ∎
@@
Theorem C.5.
Fix $10<\delta<$ 1 and $11/2<a<$ 1. Let $C\geq 128$ be a large enough universal constant for which assumption ((A1)) holds. If $n$ is suffcient large such that
@@
Notice that we only need simple concentration not uniform concentration in this theorem. We use the same definition of term $(I)$, $(II)$ as in the proof of Theorem C.4. First, by Theorem C.2, we have $(II)\leq 1/2$. Since $0\leq w_{i}(X)\leq 1$, by a standard concentration of bounded variables, we can get
@@
By Theorem C.2, we have $(VI)\leq 1/2$. For $(V)$, by triangle inequality,
@@
Using Lemma B.1 and Lemma B.2 in [50], we can get $\|w_{i}(X_{\ell})X_{\ell}\|_{\psi_{2}}\leq\|X_{\ell}\|_{\psi_{2}}\leq\tilde{c}_{3}R_{max}$, $\forall i\in[K]$. Hence by Lemma B.1 in [40], with probability at least $1-\delta_{2}$,
@@
where $\tilde{c}_{4}$ is an universal constant. And by Theorem C.2, we have
@@
Combining all terms together and taking $\delta_{1}=\delta_{2}=\delta/2$ we can bound Eq. 16 by
@@
C.3.3 Convergence results for transformer-based EM in Section C.2
@@
Remark C.6.
There is a small improvement $M=\lceil 2\log A/\delta\rceil+1$ compared to $M=\lceil 2A/\delta\rceil+1$ in [33]. Further more, it is easy to check that $\log_{\delta}(x)\leq-\log A$ for $x\in[0,1/A]$.
Lemma C.5.
@@
we have $\phi_{\delta}$ is non-decreasing on $[-A,A]$, and
@@
Similar to Lemma C.4. Omitted. ∎
Lemma C.6 (Lemma A.1 in [4]).
@@
This follows directly from Lemma C.7 and simple calculations. ∎
Now we propose the results for transformer-based EM. Similar to Section C.2, we use notations with superscript “^” to represent the output of the transformer-based EM.
Theorem C.6.
Fix $10<\delta<$ 1 and $11/2<a<$ 1. Let $C\geq 128$ be a large enough universal constant for which assumption ((A1)) holds. If $n$ is sufficient large and $\varepsilon\leq 1/100$ sufficient small such that
$\frac{\tilde{c}_{1}}{(1-a)\pi_{min}}\sqrt{\frac{R_{max}(R_{max}\vee d)\log{\left(\frac{12K}{\delta}\right)}}{n}}+\tilde{c}_{2}\left(R_{\max}+d\left(1+\sqrt{\frac{2\log(\frac{2n}{\delta})}{d}}\right)\right)n\varepsilon<\frac{1}{2}\left(a-\frac{1}{2}\right),$
@@
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},M=O(1),\log B_{{\boldsymbol{\Theta}}}\leq O{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
@@
We first claim that with probability at least $21-\delta/$ 2,
@@
Then by Theorem C.5, with probability at least $1-\delta$, we have
@@
Now we only need to prove Eq. 19. By the construction in Section C.2, we can see that $w_{\ell i}$ in first step can be well calculated, thus $|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|=0$ and the error comes only from the calculation of ${\{\hat{\mu}_{i}^{+}\}}$. Recall that $\mu_{i}^{+}=\frac{\sum_{\ell=1}^{n}w_{\ell i}X_{\ell}}{\sum_{\ell=1}^{n}w_{\ell i}}$ and
@@
By the initial condition Eq. 10 and Eq. 9, we have
@@
[…]
@@
Combine all things together, we get that with probability at least $21-\delta/$ 2,
@@
Thus taking $A=\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)}$ and and $\delta=\varepsilon$ in Lemma C.4, we can get $|\log-\widehat{\log}|\big{|}_{[1/A,A]}\leq\varepsilon$. Then by Lemma C.7, we have
@@
Thus Eq. 19 is proved. The parameter bounds can be directly computed by the construction in Section C.2 and Lemma C.4.
Theorem C.7.
Fix $10<\delta,\beta<$ 1 and $11/2<a<$ 1. Let $C\geq 128$ be a large enough universal constant. Suppose the separation condition Eq. 8 holds and suppose the initialization parameter input to transformer satisfies ${D}_{m}\leq a$. If $n$ is suffcient large and $K_{0}\varepsilon\leq 1/100$ sufficient small such that
$\epsilon(n,\varepsilon,\delta,a):=\frac{\tilde{c}_{1}}{\left(1-a\right)\pi_{\min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}n}{\delta})}{n}}+\tilde{c}_{2}{\left(\frac{1}{\pi_{\min}}+n{\left(R_{\max}+d+\sqrt{2d\log\left(\frac{2n}{\delta}\right)}\right)}\right)}\varepsilon<a(1-\beta),$
@@
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},M=O(1),\log B_{{\boldsymbol{\Theta}}}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
@@
Similar to the proof of Theorem C.6, using Theorem C.4, we only need to prove that with probability at least $21-\delta/$ 2,
@@
Define $u_{\ell}=(u_{\ell,1},\cdots,u_{\ell,K})^{\top}$, $\hat{u}_{\ell}=(\hat{u}_{\ell,1},\cdots,\hat{u}_{\ell,K})^{\top}$, where $u_{\ell,i}=\log\pi_{i}+\mu_{i}^{\top}X_{\ell}-1/2\|\mu_{i}\|^{2}$ and $\hat{u}_{\ell,i}=\widehat{\log}\pi_{i}+\mu_{i}^{\top}X_{\ell}-1/2\widehat{\|\mu_{i}\|^{2}}$ By the construction in Section C.2 and Corollary C.1, we have

### v2

@@ F.1 Formal statement of Theorem˜1
where $\overline{X}_{i}=[{X}_{i}^{\top},{\mathbf{0}}_{d_{0}-d}^{\top}]^{\top}$, $\overline{{\boldsymbol{\pi}}}_{\log}=[{\boldsymbol{\pi}}_{\log}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{\mu}_{i\%K}=[\mu_{i\%K}^{\top}$, ${\mathbf{0}}_{d_{0}-d}^{\top}]^{\top}$, $\overline{{\mathbf{w}}}_{i}=[{\mathbf{w}}_{i}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{{\mathbf{w}}}_{i\log}=[{\mathbf{w}}_{i\log}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $\overline{{\boldsymbol{\pi}}}=[{\boldsymbol{\pi}}^{\top},{\mathbf{0}}_{K_{0}-K}^{\top}]^{\top}$, $c_{i\%K}\in\mathbb{R}$ and ${\mathbf{e}}_{i\%K}\in\mathbb{R}^{K_{0}}$ denotes the $i\%K$}-th standard unit vector. For the initialization, we choose ${\boldsymbol{\pi}}_{\log}=\log{\boldsymbol{\pi}}^{(0)}$, $\mu_{i}=\mu_{i}^{(0)}$, $c_{i}=\|\mu_{i}^{(0)}\|_{2}^{2}$. and ${{\boldsymbol{\pi}}}={\mathbf{w}}_{i}={\mathbf{w}}_{i\log}={\mathbf{0}}_{K}$, $i\in[K]$. Finally, take ${\mathbf{H}}^{(0)}={\mathbf{H}}$ which is defined in Section˜F.2.
@@
$\widetilde{{\mathbf{h}}}_{i}^{(1)}$ $={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(1)}}({\mathbf{H}}^{(0)})\right]}_{:,i}$
$={\mathbf{h}}_{i}^{(0)}+\sum_{j=1}^{N}{\left[\operatorname{\mathbf{softmax}}{\left({\left({\left({\mathbf{Q}}^{(1)}{\mathbf{h}}_{i}^{(0)}\right)}^{\top}{\left({\mathbf{K}}^{(1)}{\mathbf{h}}_{j}^{(0)}\right)}\right)}_{j=1}^{N}\right)}\right]}_{j}\cdot{\mathbf{V}}^{(1)}{\mathbf{h}}_{j}^{(0)}$
@@
where ${\overline{\hat{{\mathbf{w}}}}}_{i\log}^{(1)}=\widehat{\log}{\overline{\hat{{\mathbf{w}}}}}_{i}^{(1)}$. Notice that although $\log x$ is not defined at 0, the MLP approximation is well defined with some value which we do not care because we will not use it in the M-step. Similarly, for any $\ell\%2=1$, $\ell\in\mathbb{N}_{+}$ we have
@@
$\widetilde{{\mathbf{h}}}_{j}^{(2)}$ $={\left[{\rm Attn}_{{\boldsymbol{\Theta}}_{\tt attn}^{(2)}}({\mathbf{H}}^{(1)})\right]}_{:,j}$
$={\mathbf{h}}_{j}^{(1)}+\sum_{m=1}^{2}\sum_{i=1}^{N}{\left[\operatorname{\mathbf{softmax}}{\left({\left({\left({\mathbf{Q}}_{m}^{(2)}{\mathbf{h}}_{i}^{(1)}\right)}^{\top}{\left({\mathbf{K}}_{m}^{(2)}{\mathbf{h}}_{j}^{(1)}\right)}\right)}_{j=1}^{N}\right)}\right]}_{j}\cdot{\mathbf{V}}_{m}^{(2)}{\mathbf{h}}_{i}^{(1)}$
@@
Similarly, for any $\ell\%2=0$, $\ell\in\mathbb{N}_{+}$ we have
@@
Thus, we can get $\hat{{\boldsymbol{\pi}}}^{(\ell)}$ and $\hat{\mu}_{j}^{(\ell)}$, $j\in[K]$ after $2\ell$ layers of transformer constructed above. (The last-layer MLP block retains $\boldsymbol{\pi}$} as an output parameter without cleaning it.) Our transformer construction is summarized in Fig.˜9, which is the formal version of Fig.˜7 in Section˜4.2.
Remark F.3.
@@
F.3 Convergence results for EM algorithm
F.3.1 Convergence results for population-EM algorithm
First, we review some notations. Recall that $\pi_{min}=\min_{i}\pi_{i}^{*}$, $\rho_{\pi}=\max_{i}\pi_{i}^{*}/\min_{i}\pi_{i}^{*}$, $R_{ij}=\|\mu_{i}^{*}-\mu_{j}^{*}\|$, $R_{min}=\min_{i\neq j}R_{ij}$ and $R_{max}={\left(\max_{i\neq j}R_{ij}\right)}\vee{\left(\max_{i\in[K]}\left\|{\mu_{i}^{\ast}}\right\|\right)}$. Without the loss of generality, we assume that $R_{max}\geq 1$. For clarity, we restate assumption ((A1)), which is consistent with Kwon and Caramanis (2020).
@@ F.3.1 Convergence results for population-EM algorithm
The following results gives linear convergenve guarantees of population-EM, which comes from Kwon and Caramanis (2020).
Theorem F.2 (Kwon and Caramanis (2020), Theorem 1, part i).
@@
Theorem F.3 (Kwon and Caramanis (2020), Theorem 1, part ii).
Let $C\geq 128$ be a large enough universal constant. Fix $0<a<1$. Suppose the separation condition Eq.˜7 holds and suppose the initialization parameter satisfies $D_{m}\leq a$, then $D_{m}^{+}\leq\beta D_{m}$ for some $0<\beta<1$.
Remark F.4.
Here the contraction parameter $\beta$ is only dependent with $C$ and $a$. In other words, if we fix $a\in(0,1)$, then for any $\beta\in(0,1)$, there exists a large enough $C$ such that Theorem˜F.3 holds. For details, see Appendix E in Kwon and Caramanis (2020).
Combing Theorem˜F.2 and Theorem˜F.3, we can get the linear convergence of population-EM algorithm.
F.3.2 Convergence results for empirical-EM algorithm
Now we consider the empirical-EM, i.e., Algorithm˜B.1. For convenience, the algorithm can be presented as
@@ F.3.2 Convergence results for empirical-EM algorithm
Theorem F.4.
Fix $0<\delta,\beta<1$ and $0<a<1$. Let $C\geq 128$ be a large enough universal constant. Suppose the separation condition Eq.˜7 holds and suppose the initialization parameter satisfies $D_{m}\leq a$. If $n$ is suffcient large such that
@@
where $(i)$ follows from ${|\pi_{i}^{+}-\pi_{i}^{*}|}/{\pi_{i}^{*}}\leq a$. And by Lemma˜F.3, we have
@@
In conclusion, if we take $\tilde{\beta}=(1-a)/(1+a)\beta$, $\tilde{c}=\tilde{c}_{1}\vee\tilde{c}_{2}$, $C=C(\beta,a)\geq 128$ large enough such that Theorem˜F.3 holds, and take $\delta_{1}=\delta_{2}=\delta/2$ and use union bound argument, then we get $D_{m}^{+}\leq\beta D_{m}+\varepsilon_{unif}\leq a$.
@@
Lemma F.1 (Segol and Nadler (2021), Lemma B.2.).
Fix $0<\delta<1$. Let $B_{1},\ldots,B_{K}\subset\mathbb{R}^{d}$ be Euclidean balls of radii $r_{1},\ldots,r_{K}$. Define $\mathcal{B}=\otimes_{k=1}^{K}B_{k}\subset\mathbb{R}^{Kd}$ and $r=\max_{k\in[K]}r_{k}$. Let $X$ be a random vector in $\mathbb{R}^{d}$ and $W:\mathbb{R}^{d}\times{\cal B}\to\mathbb{R}^{k}$ where $k\leq d$. Assume the following hold:
@@
Remark F.5.
There is one difference between Lemma˜F.1 and LemmaB.2. in Segol and Nadler (2021): in Lemma˜F.1, we use $1+\frac{12nLr}{\delta}$ to replace $\frac{18nLr}{\delta}$, thus we avoid the condition $r_{1},\cdots,r_{K}\geq 1$.
Hence we can get the uniform convergence of $w_{i}(X,\theta)$ and $w_{i}(X,\theta)(X-\mu_{i}^{*})$, $i\in[K]$. Our proof is similar to Segol and Nadler (2021), except that we consider the variation of both $\pi$ and $\mu$. From now on, we denote $\theta_{i}=\{\pi_{i},\mu_{i}\}$, $\theta=\{\theta_{i}\}_{i=1}^{n}$.
Lemma F.2.
Fix $0<\delta<1$ and $0<a<1$. Consider the parameter region ${\mathcal{D}}_{a}:=\{D_{m}\leq a\}$. Let $X_{1},\cdots,X_{n}\overset{\text{i.i.d.}}{\sim}\text{GMM}(\pi^{*},\mu^{*})$, then with probability at least $1-\delta$,
@@
The proof is similar to the proof of Lemma 5.1 in Segol and Nadler (2021). For simplicity, we omit it. ∎
Lemma F.3.
Fix $0<\delta<1$ and $0<a<1$. Consider the parameter region ${\mathcal{D}}_{a}:=\{D_{m}\leq a\}$. Let $X_{1},\cdots,X_{n}\overset{\text{i.i.d.}}{\sim}\text{GMM}(\pi^{*},\mu^{*})$ with $R_{min}$ satisfying Eq.˜7, then with probability at least $1-\delta$,
@@
The proof is similar to the proof of Lemma 5.4 in Segol and Nadler (2021)(Notice that the condition (36) in Segol and Nadler (2021) is trivial in our case). For simplicity, we omit it. ∎
@@
Theorem F.5.
Fix $0<\delta<1$ and $1/2<a<1$. Let $C\geq 128$ be a large enough universal constant for which assumption ((A1)) holds. If $n$ is suffcient large such that
@@
Notice that we only need simple concentration not uniform concentration in this theorem. We use the same definition of term $(I)$, $(II)$ as in the proof of Theorem˜F.4. First, by Theorem˜F.2, we have $(II)\leq 1/2$. Since $0\leq w_{i}(X)\leq 1$, by a standard concentration of bounded variables, we can get
@@
By Theorem˜F.2, we have $(VI)\leq 1/2$. For $(V)$, by triangle inequality,
@@
Using Lemma B.1 and Lemma B.2 in Zhao et al. (2020), we can get $\|w_{i}(X_{\ell})X_{\ell}\|_{\psi_{2}}\leq\|X_{\ell}\|_{\psi_{2}}\leq\tilde{c}_{3}R_{max}$, $\forall i\in[K]$. Hence by Lemma B.1 in Segol and Nadler (2021), with probability at least $1-\delta_{2}$,
@@
where $\tilde{c}_{4}$ is an universal constant. And by Theorem˜F.2, we have
@@
Combining all terms together and taking $\delta_{1}=\delta_{2}=\delta/2$ we can bound Eq.˜15 by
@@
F.3.3 Convergence results for transformer-based EM in Section˜F.2
@@
Remark F.6.
There is a small improvement $M=\lceil 2\log A/\delta\rceil+1$ compared to $M=\lceil 2A/\delta\rceil+1$ in Mei (2024). Further more, it is easy to check that $\log_{\delta}(x)\leq-\log A$ for $x\in[0,1/A]$.
Lemma F.5.
@@
we have $\phi_{\delta}$ is non-decreasing on [A,A][-A,A], and
@@
Similar to Lemma˜F.4. Omitted. ∎
Lemma F.6 (Lemma A.1 in Bai et al. (2023)).
@@
This follows directly from Lemma˜F.7 and simple calculations. ∎
Now we propose the results for transformer-based EM. Similar to Section˜F.2, we use notations with superscript “ˆ” to represent the output of the transformer-based EM.
Theorem F.6.
Fix $0<\delta<1$ and $1/2<a<1$. Let $C\geq 128$ be a large enough universal constant for which assumption ((A1)) holds. If $n$ is sufficient large and $\varepsilon\leq 1/100$ sufficient small such that
$\frac{\tilde{c}_{1}}{(1-a)\pi_{min}}\sqrt{\frac{R_{max}(R_{max}\vee d)\log{\left(\frac{12K}{\delta}\right)}}{n}}$
$\qquad+\tilde{c}_{2}\left(R_{\max}+d\left(1+\sqrt{\frac{2\log(\frac{2n}{\delta})}{d}}\right)\right)n\varepsilon<\frac{1}{2}\left(a-\frac{1}{2}\right),$
@@
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},$
$M=O(1),$
$\log B_{{\boldsymbol{\Theta}}}\leq O{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
@@
We first claim that with probability at least $1-\delta/2$,
@@
Then by Theorem˜F.5, with probability at least $1-\delta$, we have
@@
Now we only need to prove Eq.˜18. By the construction in Section˜F.2, we can see that $w_{\ell i}$ in first step can be well calculated, thus $|\hat{\pi}_{i}^{+}-\pi_{i}^{+}|=0$ and the error comes only from the calculation of $\{\hat{\mu}_{i}^{+}\}$}. Recall that $\mu_{i}^{+}=\frac{\sum_{\ell=1}^{n}w_{\ell i}X_{\ell}}{\sum_{\ell=1}^{n}w_{\ell i}}$ and
@@
By the initial condition Eq.˜9 and Eq.˜8, we have
@@
Since $X_{\ell}\overset{\text{i.i.d.}}{\sim}\text{GMM}(\pi^{*},\mu^{*})$, using Lemma˜F.6, with probability at least $1-\delta/2$, we have
@@
Combine all things together, we get that with probability at least $1-\delta/2$,
@@
Thus taking $A=\exp{\left(\tilde{O}(K_{0}R_{\max}(R_{\max}+d_{0}))\right)}$ and and $\delta=\varepsilon$ in Lemma˜F.4, we can get $|\log-\widehat{\log}|\big|_{[1/A,A]}\leq\varepsilon$. Then by Lemma˜F.7, we have
@@
Thus Eq.˜18 is proved. The parameter bounds can be directly computed by the construction in Section˜F.2 and Lemma˜F.4.
Theorem F.7.
Fix $0<\delta,\beta<1$ and $1/2<a<1$. Let $C\geq 128$ be a large enough universal constant. Suppose the separation condition Eq.˜7 holds and suppose the initialization parameter input to transformer satisfies ${D}_{m}\leq a$. If $n$ is suffcient large and $K_{0}\varepsilon\leq 1/100$ sufficient small such that
$\epsilon(n,\varepsilon,\delta,a)$ $:=\frac{\tilde{c}_{1}}{\left(1-a\right)\pi_{\min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}n}{\delta})}{n}}+\tilde{c}_{2}{\left(\frac{1}{\pi_{\min}}+n{\left(R_{\max}+d+\sqrt{2d\log\left(\frac{2n}{\delta}\right)}\right)}\right)}\varepsilon$
$<a(1-\beta),$
@@
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},$
$M=O(1),$
$\log B_{{\boldsymbol{\Theta}}}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
@@
Similar to the proof of Theorem˜F.6, using Theorem˜F.4, we only need to prove that with probability at least $1-\delta/2$,
@@
Define $u_{\ell}=(u_{\ell,1},\cdots,u_{\ell,K})^{\top}$, $\hat{u}_{\ell}=(\hat{u}_{\ell,1},\cdots,\hat{u}_{\ell,K})^{\top}$, where $u_{\ell,i}=\log\pi_{i}+\mu_{i}^{\top}X_{\ell}-1/2\|\mu_{i}\|^{2}$ and $\hat{u}_{\ell,i}=\widehat{\log}\pi_{i}+\mu_{i}^{\top}X_{\ell}-1/2\widehat{\|\mu_{i}\|^{2}}$ By the construction in Section˜F.2 and Corollary˜F.1, we have